# Finding Kissing Numbers with Game-theoretic Reinforcement Learning

### v1

@@ Abstract
Since Isaac Newton first studied the Kissing Number Problem in 1694, determining the maximal number of non-overlapping spheres around a central sphere has remained a fundamental challenge. This problem represents the local analogue of Hilbert’s 18th problem on sphere packing, bridging geometry, number theory, and information theory. Although significant progress has been made through lattices and codes, the irregularities of high-dimensional geometry and exponentially growing combinatorial complexity beyond 8 dimensions, which exceeds the complexity of Go game, limit the scalability of existing methods. Here we model this problem as a two-player matrix completion game and train the game-theoretic reinforcement learning system, PackingStar, to efficiently explore high-dimensional spaces. The matrix entries represent pairwise cosines of sphere center vectors; one player fills entries while another corrects suboptimal ones, jointly maximizing the matrix size, corresponding to the kissing number. This cooperative dynamics substantially improves sample quality, making the extremely large spaces tractable. PackingStar reproduces previous configurations and surpasses all human-known records from dimensions 25 to 31, with the configuration in 25 dimensions geometrically corresponding to the Leech lattice and suggesting possible optimality. It achieves the first breakthrough beyond rational structures from 1971 in 13 dimensions and discovers over 6000 new structures in 14 and other dimensions. These results demonstrate AI’s power to explore high-dimensional spaces beyond human intuition and open new pathways for the Kissing Number Problem and broader geometry problems.
@@ Introduction
PackingStar demonstrates remarkable scalability, successfully reproducing all known kissing number configurations in dimensions up to 31. Figure 1 illustrates some of PackingStar’s breakthrough discoveries on this problem. From dimensions 25 to 31, PackingStar surpasses the lower bounds reported in 2016 and also exceeds the meta-construction form introduced in 2011 [?,?], achieving what appears to be the optimal construction form along this path; the resulting construction form are summarized in Table 2. Notably, the breakthrough configuration discovered in 25 dimensions exhibits a clear geometric pattern that corresponds closely to the substructure of Leech lattice. Across multiple metrics, this configuration shows strong convergence, suggesting that it may represent the optimal 25-dimensional arrangement, although a rigorous mathematical proof is still lacking. In 13 dimensions, PackingStar surpasses the previously known rational structure with 1130 spheres from 1971 [?], reaching a kissing number of 1146. Rational structures are those in which all pairwise cosines between sphere centers are rational. The non-rational structure with 1154 spheres is obtained by slightly adjusting the coordinates of some spheres in the 1130-sphere configuration, later reached 1154 in 1999 [?]. Although the new 13-dimensional rational configuration does not exceed the non-rational record, it is of considerable mathematical value. Its fully rational inter-sphere cosines allow for exact analytic study without relying on numerical approximations. This offers insights into the effects of different constraints on kissing numbers and may inspire higher-density packings. Rational structures also facilitate more efficient encoding and decoding in information and communication applications, highlighting their practical significance alongside their theoretical and geometric value. Beyond these specific dimensions, PackingStar uncovers over 6000 new configurations in 14 dimensions and identifies additional configurations corresponding to current lower bounds in dimensions 12, 13, 15, and others. The sheer volume of these newly discovered arrangements provides rich opportunities for further investigation and may ultimately lead to configurations with even higher kissing numbers. Collectively, these results highlight AI’s ability to understand and reason about high-dimensional geometric structures, opening a new paradigm for studying the Kissing Number Problem and demonstrating the potential of AI to explore a broader class of challenging mathematical problems.
@@ Results
We evaluate the PackingStar system across multiple high-dimensional regimes and report new state-of-the-art kissing configurations. Our results fall into three categories: (i) new lower bounds in dimensions 25253131, (ii) improved rational constructions in $13$ dimensions, and (iii) diverse families of solutions in different dimensions. These findings demonstrate that PackingStar can both recover classical lattice-based structures and uncover entirely new high-dimensional arrangements, substantially expanding the landscape of known kissing configurations. The main kissing configurations are available at the data repository https://github.com/CDM1619/PackingStar.
@@
Table 2: New kissing numbers $K_{\mathrm{new}}(n)$ by PackingStar vs. previous $K_{\mathrm{prev}}(n)$ for $25\leq n\leq 31$. The asterisk $\bigstar$ indicates that under our new construction form, the theoretical lower bound in 31 dimensions should be 238350. The present version reports 238078 only due to computational and time limitations, rather than any intrinsic limitation of our method. The Future versions are expected to reach the theoretical bound. In the new construction form, bold items indicate the improved configuration forms discovered by PackingStar compared with previous forms.
@@
In this setting, the first objective is to identify the largest possible subsets $S_{i}$. Each $S_{i}$ corresponds to a subset of shortest Leech lattice vectors whose pairwise cosine values are bounded by $1/4$; equivalently, in the PackingStar framework this corresponds to cosine sets $C_{1}=C_{2}=\{-1,\ 0,\ \pm 1/4\}$ together with the structural constraint that $C_{*}$ matches the Leech minimal-vector set. Previous work identified subsets $|S_{i}|=488$ with no apparent geometric structure [?]. PackingStar discovers subsets $|S_{i}|=496$ exhibiting a clear internal geometric organization. Specifically, this 496-element subset is composed of 28 symmetric-frame structures $\mathrm{X}_{8}$ in 8-dimensional subspaces, each containing 16 spheres, together with a symmetric-frame structure (“Conway-Curtis Cross” [?]) $\mathrm{X}_{24}$ in 24-dimensional space containing 48 spheres, which corresponds to an important substructure of the Leech lattice. Here the symmetric-frame structure is a Cross-polytope. Based on this clear geometric interpretation, we conjecture that this subset $S_{i}$ may be optimal, and consequently, the 25-dimensional kissing configuration derived from this construction could also attain the maximum kissing number.
@@
The second objective is to identify improved meta construction forms that assemble these subsets into full configurations. Both objectives are jointly addressed within the unified PackingStar framework, which simultaneously increases the achievable size of $S_{i}$ and uncovers superior meta construction forms. Our findings show that PackingStar not only discovers larger subsets but also identifies a strictly better construction template than the meta-construction introduced in 2011; moreover, evidence from convergence and geometric consistency strongly suggests that the new template is optimal within this construction pathway. These advances yield especially pronounced improvements in dimensions 29 and 31. Table 2 reports additional new construction forms that deviate from the classical form [?,?]. In dimensions 26263030, the system consistently augments the Leech-lattice component by adding lower-dimensional parts $K(k)$ where k=n24k=n-24 (e.g., adding $K(2)$ in dimension $26$). In 31 dimensions, PackingStar uncovers a novel assembly pattern by partitioning a 7-dimensional kissing configuration into 42 (optimal) disjoint unit-radius equilateral triangles, resulting in an $84$-fold weighted $S_{i}$, surpassing the previous $75$-fold weighted $S_{i}$. Similarly, in 29 dimensions, PackingStar embeds 12 (optimal) disjoint unit-radius equilateral triangles into the 5-dimensional kissing configuration to generate a $26$-fold weighted $S_{i}$, exceeding the prior $24$-fold weighted $S_{i}$. These advances yield strictly larger kissing configurations in all dimensions from $25$ to $31$. The resulting new lower bounds $K_{\mathrm{new}}(n)$ are summarized in Table 2. Compared with the previous best values $K_{\mathrm{prev}}(n)$, our constructions increase the kissing number by $8,38,68,152,1224,456,$ and $5204$ spheres in dimensions $25$ through $31$, respectively. These improvements are also highlighted by the red markers in Fig. 1. Table 3 summarizes the cosine value among sphere pairs in the new configurations.
@@ Methods
In this section, we present a learning-based system PackingStar to find high-dimensional kissing constructions. First, we extract dominant geometric features of candidate kissing configurations via simulation and Monte Carlo Tree Search. Second, we formulate the search for maximal kissing numbers as a cooperative two-player matrix completion game, where one agent iteratively fills the Gram matrix and the other corrects suboptimal entries, jointly maximizing the final matrix size.
@@
At each step $m\geq m_{0}$, Player 1 samples an action from policy $\pi_{1}$, $\mathbf{g}\sim\pi_{1}(\cdot\mid\mathbf{G}^{(m)})$, where $\mathbf{g}\in\mathcal{A}^{(m)}$ to extend the current matrix $\mathbf{G}^{(m)}\in\mathbb{R}^{m\times m}$:
@@
$\mathbf{g}=\begin{bmatrix}\mathbf{g}^{(1)}\\ \mathbf{g}^{(2)}\end{bmatrix},\quad\mathbf{g}^{(1)}\in\mathbb{R}^{n},\ \mathbf{g}^{(2)}\in\mathbb{R}^{m-n}.$ (6)
@@
$\mathbf{g}^{(2)}={\mathbf{G}^{\prime}}(\mathbf{G}^{(m)}_{1:n,\,1:n})^{+}\mathbf{g}^{(1)}\,.$ (8)
@@
$I\sim\pi_{2}(\cdot\mid\mathbf{G}^{(M)})\subseteq\{1,2,\dots,M\},$ (12)
and constructs the refined matrix
@@
$\mathbf{G}^{(M)}\xrightarrow{\pi_{2}}\mathbf{G}^{*}\xrightarrow{\pi_{1}}\mathbf{G}^{(M^{\prime})}\xrightarrow{\pi_{2}}\cdots$ (14)
@@ Conclusion
Trained from scratch without any human-designed priors, PackingStar automatically discovers high-dimensional kissing configurations that outperform all known constructions and exhibit clear mathematical structure. The key idea is to reformulate the Kissing Number Problem as a two-player matrix completion game over the Gram-matrix domain, enabling efficient exploration of an extremely large and sparse space without explicit coordinate representations. Notably, PackingStar improves all known lower bounds from dimensions 25 to 31. In 13 dimensions, it uncovers a rational configuration with a larger kissing number; and in 14 and other dimensions, it identifies thousands of new configurations that provide a rich experimental basis for structural classification and better geometric analysis. A current limitation is that PackingStar cannot provide optimality proofs for these new configurations. We anticipate that the interaction between reinforcement learning and rigorous proof techniques will not only open new pathways for the Kissing Number Problem, but also extend to broader geometry problems.
@@ Acknowledgments
We appreciate Ao Li’s assistance in program optimization and large-scale computing cluster deployment. We thank Mingzhi Wang and Juntao Dai for their assistance in coordinating GPU resources. We are grateful to Henry Cohn for his influential work on the Kissing Number Problem, which has been a source of inspiration for this research. We thank Chiyuan Wang and Xinmian Sun for the early discussions on this topic.
@@
C.M. conceived and led the project. Z.T. established geometric metrics and analyzed the properties of the structures. C.M. developed the game-theoretic reinforcement learning method. M.L., Z.T., Z.M. and C.M. explored geometric relationships between structures of different dimensions. H.C. explored the training of geometric generative models for high-dimensional structures. C.M., H.C., and M.L. modeled the process of cosine feature simulation and designed the simulation algorithm. M.L., C.M. and Z.T. discovered the cosine feature of constructions via numerical simulation. C.M., H.C., Z.T., M.L., P.L., Z.M. and Y.Y. completed the formulation of matrix completion and designed the algorithm. Z.M. provided practical insights and metrics on the exploration of lattices and error-correcting codes, particularly in the construction of 31D, which provides core insights. C.M., P.L., M.L., Z.T. and H.C. developed the project code. C.M., Z.T., P.L. and H.C. conducted the experiments. C.M. and Y.Y. wrote the manuscript. Y.Q., Y.Y. and Y.C. provided computational resources and contributed to the discussions.

### v2

@@ Abstract
Since Isaac Newton first studied the Kissing Number Problem in 1694, determining the maximal number of non-overlapping spheres around a central sphere has remained a fundamental challenge. This problem represents the local analogue of Hilbert’s 18th problem on sphere packing, bridging geometry, number theory, and information theory. Although significant progress has been made through lattices and codes, the irregularities of high-dimensional geometry and exponentially growing combinatorial complexity beyond 8 dimensions, which exceeds the complexity of Go game, limit the scalability of existing methods. Here we model this problem as a two-player matrix completion game that can be fully parallelized at large scale, and train the game-theoretic reinforcement learning system, PackingStar, to efficiently explore high-dimensional spaces. The matrix entries represent pairwise cosines of sphere center vectors; one player fills entries while another corrects suboptimal ones, jointly maximizing the matrix size, corresponding to the kissing number. This cooperative dynamics substantially improves sample quality, making the extremely large spaces tractable. PackingStar reproduces previous configurations and surpasses all human-known records from dimensions 25 to 31, with the configuration in 25 dimensions geometrically corresponding to the Leech lattice and suggesting possible optimality. It achieves the first breakthrough beyond rational structures from 1971 in 13 dimensions, discovers over 6000 new structures in 14 and other dimensions, and sets new records for generalized kissing configurations under various angular constraints. These results demonstrate AI’s power to explore high-dimensional spaces beyond human intuition and open new pathways for the Kissing Number Problem and broader geometry problems.
@@ Introduction
PackingStar demonstrates remarkable scalability, successfully reproducing all known kissing number configurations in dimensions up to 31. Figure 1 illustrates some of PackingStar’s breakthrough discoveries on this problem. From dimensions 25 to 31, PackingStar surpasses the lower bounds reported in 2016 and also exceeds the meta-construction form introduced in 2011 [?,?], achieving what appears to be the optimal construction form along this path; the resulting construction form are summarized in Table 2. Notably, the breakthrough configuration discovered in 25 dimensions exhibits a clear geometric pattern that corresponds closely to the substructure of Leech lattice. Across multiple metrics, this configuration shows strong convergence, suggesting that it may represent the optimal 25-dimensional arrangement, although a rigorous mathematical proof is still lacking. In 13 dimensions, PackingStar surpasses the previously known rational structure with 1130 spheres from 1971 [?], reaching a kissing number of 1146. Rational structures are those in which all pairwise cosines between sphere centers are rational. The non-rational structure with 1154 spheres is obtained by slightly adjusting the coordinates of some spheres in the 1130-sphere configuration, later reached 1154 in 1999 [?]. Although the new 13-dimensional rational configuration does not exceed the non-rational record, it is of considerable mathematical value. Its fully rational inter-sphere cosines allow for exact analytic study without relying on numerical approximations. This offers insights into the effects of different constraints on kissing numbers and may inspire higher-density packings. Rational structures also facilitate more efficient encoding and decoding in information and communication applications, highlighting their practical significance alongside their theoretical and geometric value. Beyond these specific dimensions, PackingStar uncovers over 6000 new configurations in 14 dimensions and identifies additional configurations corresponding to current lower bounds in dimensions 12, 13, 15, and others. The sheer volume of these newly discovered arrangements provides rich opportunities for further investigation and may ultimately lead to configurations with even higher kissing numbers. In addition, PackingStar achieves significant new records for generalized kissing configurations under tighter angular constraints, including a 12-dimensional configuration with 81 spheres under the $1/4$ cosine constraint, as well as 14-dimensional and 17-dimensional configurations with 252 and 578 spheres under the $1/3$ constraint, respectively. All of these results have been independently validated and are now included in the standard databases for kissing numbers and spherical codes. In summary, these results highlight AI’s ability to understand and reason about high-dimensional geometric structures, opening a new paradigm for studying the Kissing Number Problem and demonstrating the potential of AI to explore a broader class of challenging mathematical problems.
@@
In summary, AlphaEvolve demonstrates that AI can “optimize” geometry, PackingStar shows that AI can truly “create” geometry. This represents the first genuinely systematic breakthrough of AI in high-dimensional discrete geometry and coding theory: it is not merely a numerical improvement, but a methodological advance, illustrating how AI can uncover new structures and insights in previously unexplored high-dimensional spaces.
@@ Results
We evaluate the PackingStar system across multiple high-dimensional regimes and report new state-of-the-art kissing configurations. Our results fall into three categories: (i) new lower bounds in dimensions 25253131, (ii) improved rational constructions in $13$ dimensions, (iii) diverse families of solutions in different dimensions, and (iv) new records for generalized kissing configurations. These findings demonstrate that PackingStar can both recover classical lattice-based structures and uncover entirely new high-dimensional arrangements, substantially expanding the landscape of known kissing configurations. The main kissing configurations are available at the data repository https://github.com/CDM1619/PackingStar.
@@
Table 2: New kissing numbers $K_{\mathrm{new}}(n)$ by PackingStar vs. previous $K_{\mathrm{prev}}(n)$ for $25\leq n\leq 31$. In the new construction form, bold items indicate the improved configuration forms discovered by PackingStar compared with previous forms [?].
@@
In this setting, the first objective is to identify the largest possible subsets $S_{i}$. Each $S_{i}$ corresponds to a subset of shortest Leech lattice vectors whose pairwise cosine values are bounded by $1/4$; equivalently, in the PackingStar framework this corresponds to cosine sets $C_{1}=C_{2}=\{-1,\ 0,\ \pm 1/4\}$ together with the structural constraint that $C_{*}$ matches the Leech lattice minimal-vector set. Previous work identified subsets $|S_{i}|=488$ with no apparent geometric structure [?]. PackingStar discovers subsets $|S_{i}|=496$ exhibiting a clear internal geometric organization. Specifically, this 496-element subset is composed of 28 symmetric-frame structures $\mathrm{X}_{8}$ in 8-dimensional subspaces, each containing 16 spheres, together with a symmetric-frame structure (“Conway-Curtis Cross” [?]) $\mathrm{X}_{24}$ in 24-dimensional space containing 48 spheres, which corresponds to an important substructure of the Leech lattice. Here the symmetric-frame structure is a Cross-polytope. Based on this clear geometric interpretation, we conjecture that this subset $S_{i}$ may be optimal, and consequently, the 25-dimensional kissing configuration derived from this construction could also attain the maximum kissing number.
@@
The second objective is to identify improved meta construction forms that assemble these subsets into full configurations. Both objectives are jointly addressed within the unified PackingStar framework, which simultaneously increases the achievable size of $S_{i}$ and uncovers superior meta construction forms. Our findings show that PackingStar not only discovers larger subsets but also identifies a strictly better construction template than the meta-construction introduced in 2011; moreover, evidence from convergence and geometric consistency strongly suggests that the new template is optimal within this construction pathway. These advances yield especially pronounced improvements in dimensions 29 and 31. Table 2 reports additional new construction forms that deviate from the classical form [?,?]. In dimensions 26263030, the system consistently augments the Leech-lattice component by adding lower-dimensional parts $K(k)$ where k=n24k=n-24 (e.g., adding $K(2)$ in dimension $26$). In 31 dimensions, PackingStar uncovers a novel assembly pattern by partitioning a 7-dimensional kissing configuration into 42 (optimal) disjoint unit-radius equilateral triangles, resulting in an $84$-fold weighted $S_{i}$, surpassing the previous $75$-fold weighted $S_{i}$. Similarly, in 29 dimensions, PackingStar embeds 12 (optimal) disjoint unit-radius equilateral triangles into the 5-dimensional kissing configuration to generate a $26$-fold weighted $S_{i}$, exceeding the prior $24$-fold weighted $S_{i}$. These advances yield strictly larger kissing configurations in all dimensions from $25$ to $31$. The resulting new lower bounds $K_{\mathrm{new}}(n)$ are summarized in Table 2. Compared with the previous best values $K_{\mathrm{prev}}(n)$, our constructions increase the kissing number by $8,38,68,152,1224,456,$ and $5476$ spheres in dimensions $25$ through $31$, respectively. These improvements are also highlighted by the red markers in Fig. 1. Table 3 summarizes the cosine value among sphere pairs in the new configurations.
@@
New records for generalized kissing configurations
In this section, we apply PackingStar to explore generalized kissing configurations with maximal inner product at most $1/3$ and $1/4$. For geometric intuition, these cases may be viewed as double-sphere and triple-sphere kissing configurations, in which one asks how many additional unit spheres can be simultaneously tangent to two or three mutually tangent spheres, respectively. These configurations arise naturally as fundamental building blocks in sphere packings and play an important role in the study of the broader packing problem.
The classical kissing number is denoted by $K(n)$, which corresponds to the case in which the pairwise angles between sphere centers have cosine at most $1/2$. More generally, for an angular constraint $1<\alpha<1$ with $\alpha\neq 1/2$, we use $K(n,\alpha)$ to denote maximum number of unit vectors such that the inner product between any two distinct vectors is at most $\alpha$.
Table 4: New maximal values $K_{\mathrm{new}}(n,\alpha)$ for generalized kissing configurations with maximal inner product at most $\alpha$, obtained by PackingStar, compared with the previous best known values $K_{\mathrm{prev}}(n,\alpha)$. Bold entries indicate improved configurations discovered by PackingStar.
$n$ $\alpha$ New cosine set $\boldsymbol{K_{\mathrm{new}}(n,\alpha)}$ $K_{\mathrm{prev}}(n,\alpha)$
12 $1/4$ $\{-\tfrac{1}{2},\ -\tfrac{1}{8},\ \tfrac{1}{4}\}$ 81 79
14 $1/3$ $\{-1,\ 0,\ \pm\tfrac{1}{3}\}$; variants: $\{-1,\ -\tfrac{2}{3},\ 0,\ \pm\tfrac{1}{3}\}$ 252 240 [?]
17 $1/3$ $\{-1,\ -\tfrac{7}{9},\ -\tfrac{5}{9},\ 0,\ \pm\tfrac{1}{9},\ \pm\tfrac{1}{3}\}$ 578 546 [?]
As shown in Table 4, the new constructions from PackingStar yield new record for $K(n,\alpha)$ in several dimensions. In 12 dimensions under the $1/4$ cosine constraint, we achieve $K_{\mathrm{new}}(12,1/4)=81$, improving over the previous best of 79. Similarly, in 14 and 17 dimensions under the $1/3$ constraint, we obtain $K_{\mathrm{new}}(14,1/3)=252$ and $K_{\mathrm{new}}(17,1/3)=578$, surpassing the previously reported values of 240 and 546, respectively.
Notably, the $K_{\mathrm{new}}(12,1/4)$ configuration contains no antipodal pairs; however, each of the 81 spheres has the same number of neighbors at each distance, forming a highly regular local structure. Such a combination of high local regularity without any antipodal pair appears to be rare in spherical code. We find several distinct variants of $K_{\mathrm{new}}(12,1/4)$ configuration, among which the most symmetric realization turns out to be the Kronecker product of the Schläfli spherical code with the unit equilateral triangle. This arrangement possesses a transitive symmetry group of order 311040.
In dimension 17, PackingStar found a significant improvement: a configuration with 578 spheres, exceeding the previously known 546. All three geometrically non-isomorphic variants discovered share the same cosine set. Analysis of these variants reveals a remarkable structural property: the $K_{\mathrm{new}}(17,1/3)$ configuration can be interpreted as essentially two copies of the 16-dimensional 288-sphere configuration, which is known to be optimal under the cosine $1/4$ constraint, together with two additional points, i.e., $578=288\times 2+2$. This decomposition provides strong evidence that 578 is very likely optimal for the cosine $1/3$ constraint in dimension 17.
@@ Methods
In this section, we present a learning-based system PackingStar to find high-dimensional kissing constructions. First, we extract dominant geometric features of candidate kissing configurations via simulation and Monte Carlo Tree Search. Second, we formulate the search for maximal kissing numbers as a cooperative two-player matrix completion game, where one agent iteratively fills the Gram matrix and the other corrects suboptimal entries, jointly maximizing the final matrix size. This formulation allows all operations to be parallelized at large scale on NVIDIA GPUs using low-level mathematical operators in cuBLAS library, fully exploiting computational power to accelerate both the game simulation and model training.
@@
For each step $m\geq m_{0}$, Player 1 samples an action according to policy $\pi_{1}$, $\mathbf{g}\sim\pi_{1}(\cdot\mid\mathbf{G}^{(m)})$, where $\mathbf{g}\in\mathcal{A}^{(m)}$ to extend the current matrix $\mathbf{G}^{(m)}\in\mathbb{R}^{m\times m}$:
@@
$\mathbf{g}=\begin{bmatrix}\mathbf{g}^{(1)}\\ \mathbf{g}^{(2)}\end{bmatrix},\quad\mathbf{g}^{(1)}\in\mathbb{R}^{n\times 1},\ \mathbf{g}^{(2)}\in\mathbb{R}^{(m-n)\times 1}.$ (6)
@@
$\mathbf{g}^{(2)}={\mathbf{G}^{\prime}}(\mathbf{G}^{(m)}_{1:n,\,1:n})^{+}\mathbf{g}^{(1)}\,$ (8)
@@
$I\sim\pi_{2}(\cdot\mid\mathbf{G}^{(M)})$ (12)
where $\pi_{2}(\cdot\mid\mathbf{G}^{(M)})$ denotes a probability distribution on $\{1,2,\dots,M\}$, and constructs the refined matrix
@@
$\mathbf{G}^{(M)}\xrightarrow{\pi_{2}}\mathbf{G}^{*}\xrightarrow{\pi_{1}}\mathbf{G}^{(M^{\prime})}\xrightarrow{\pi_{2}}\cdots.$ (14)
@@ Conclusion
Trained from scratch without any human-designed priors, PackingStar automatically discovers high-dimensional kissing configurations that outperform all known constructions and exhibit clear mathematical structure. The key idea is to reformulate the Kissing Number Problem as a two-player matrix completion game over the Gram-matrix domain, enabling efficient exploration of an extremely large and sparse space without explicit coordinate representations. Notably, PackingStar improves all known lower bounds from dimensions 25 to 31 and sets new records for generalized kissing configurations under various angular constraints. A current limitation is that PackingStar cannot provide optimality proofs for these new configurations. We anticipate that the interaction between reinforcement learning and rigorous proof techniques will not only open new pathways for the Kissing Number Problem, but also extend to broader geometry problems.
@@ Acknowledgments
We are grateful to Professor Henry Cohn for his influential work on the Kissing Number Problem, which has been a source of inspiration for this research. We are also deeply grateful for his guidance and valuable insights, particularly for helping us identify the new structures we discovered and providing constructive suggestions that greatly facilitated our progress in generalized kissing configurations. We appreciate Ao Li’s assistance in program optimization and large-scale computing cluster deployment. We thank Mingzhi Wang and Juntao Dai for their assistance in coordinating GPU resources. We thank Chiyuan Wang and Xinmian Sun for the early discussions on this topic.
@@
C.M. conceived and led the project. T.T.Z. established geometric metrics and analyzed the properties of the structures. C.M. established the matrix completion game and developed the game-theoretic reinforcement learning method. M.L., T.T.Z., Z.M. and C.M. explored geometric relationships between structures of different dimensions. H.C. explored the training of geometric generative models for high-dimensional structures. C.M., H.C., and M.L. modeled the process of cosine feature simulation and designed the simulation algorithm. M.L., C.M. and T.T.Z. discovered the cosine feature of constructions via numerical simulation. C.M., H.C., T.T.Z., M.L., P.L. and Z.M. completed the formulation of matrix completion and designed the algorithm. Z.M. provided practical insights and metrics for exploring lattices and error-correcting codes, particularly for the 31D construction. C.M., P.L., M.L., T.T.Z. and H.C. developed the project code. C.M., T.T.Z., P.L. and H.C. conducted the experiments. C.M. and Y.Y. wrote the manuscript. Y.Q., Y.Y. and Y.C. provided computational resources and improved the efficiency of AI systems in parallel computation.